Circular No. 6075

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

2S 1417–624 C. A. Wilson and R. B. Wilson, Marshall Space Flight Center, NASA;

M. H. Finger and S. N. Zhang, University Space Research Association; and D. Chakrabarty and T. A. Prince, California Institute of Technology, report for the Compton Observatory BATSE Team: “An outburst of the transient x-ray binary pulsar 2S 1417–624 is currently being observed. Pulsations in the energy band 20-50 keV were first detected on Aug. 26 at a frequency consistent with the discovery observation of this source (Kelley et al. 1981, Ap.J. 243, 251). On Sept. 4–5, the source was detected between 20 and 70 keV with a phase-averaged pulsed flux of 10 ± 2 mCrab and a total flux of 75 ± 15 mCrab. The pulse phases between Aug. 26 and Sept. 6 are consistent with a barycentric pulse frequency of 5.666395(5) × 10^{−2} Hz on Aug. 31.0 UT, with a frequency derivative of 2.936(9) × 10^{−11} Hz/s. A search of the 20- to 50-keV BATSE data from 1991 April to date revealed no previous outbursts. Optical observations of the proposed Be-star companion (Grindlay et al. 1984, Ap.J. 276, 621) are encouraged.”

X-RAY NOVA IN SCORPIUS W. S. Paciesas, S. N. Zhang, C. A. Wilson, and B. A. Harmon re-

port for the Compton Gamma Ray Observatory BATSE team: “We have detected a second outburst of this object (GRO J1655–40), beginning on Sept. 6. The hard x-ray (20-100 keV) flux from this source had declined rapidly from ∼ 700 to ∼ 150 mCrab over a few-day period beginning ∼ Aug. 12 and remained at or below that level through Sept. 5. Our preliminary flux estimates for Sept. 6 and 7 are 440 ± 50 and 370 ± 30 mCrab, respectively. The rapid decline of the first x-ray outburst coincides well with the calculated emission time of the expanding radio lobe (IAUC 6073). The reflare may portend further interesting radio and/or optical activity.”

PERIODIC COMET MACHHOLZ 2 (1994o) Further total visual magnitude estimates: Sept. 2.36 UT, 7.0 (J. E.

Bortle, Stormville, NY, 10×50 binoculars); 5.12, 6.8 (H. Mikuž, Cres, Croatia, 7×50 binoculars; faint tail ∼ 1^{o} long in p.a. 310^{o}); 6.11, 7.2 (Mikuž); 7.48, 7.3 (C. S. Morris, near Mt. Wilson, CA, 20×80 binoculars).

1994 September 9 Daniel W. E. Green